OpenStudy (anonymous):

Fill in the blanks. › When graphing a system of inequalities, the graph is shaded below the line for the ____ and ____ symbols. › The solution to a system of inequalities is represented on the graph where ______________ ›

OpenStudy (anonymous):

????

OpenStudy (anonymous):

\(<,>\) for the first one i bet

OpenStudy (anonymous):

\(\leq, \geq\) for the second one sound right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

third one is a bit confusing though i am going to make a guess that they want \(>, \geq\)

OpenStudy (anonymous):

and < < - on the next one?

OpenStudy (anonymous):

what is the last blank